OpenStudy (anonymous):

convert 210 degrees to radians

OpenStudy (anonymous):

3.66519143 radians

OpenStudy (anonymous):

210*pi/180 = 7pi/6

OpenStudy (anonymous):

an easy formula is just \[\frac{ x }{ 1 } * \frac{ \pi }{ 180 } = answer\]

OpenStudy (anonymous):

radians = (degrees) * (pi/180) = (degrees) * (2pi/360)